# On the Diameter of Arrangements of Topological Disks

### v1

@@ Abstract
Let $\mathcal{D}=\{D_{0},\ldots,D_{n-1}\}$ be a set of $n$ topological disks in the plane and let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$. For two disks $D_{i},D_{j}\in\mathcal{D}$, let $\Delta_{ij}$ be the number of connected components of $D_{i}\cap D_{j}$, and let $\Delta:=\max_{i,j}\Delta_{ij}$.
We show that the diameter of $\mathcal{G}^{*}$, the dual graph of $\mathcal{A}$, can be bounded as a function of $n$ and $\Delta$. Thus, any two points in the plane can be connected by a Jordan curve that crosses the disk boundaries a number of times bounded by a function of $n$ and $\Delta$. In particular, for the case of two disks we prove that the diameter of $\mathcal{G}^{*}$ is at most $\max\{2,2\Delta\}$ and this bound is tight. For the general case of $n>2$ disks, we show that the diameter of $\mathcal{G}^{*}$ is at most $2n(\Delta+1)^{n(n-1)/2}\min\{n,\Delta+1\}$. We achieve this by proving that the number of maximal faces in $\mathcal{A}$ — the faces whose ply is more than the ply of their neighboring faces — is at most $n(\Delta+1)^{n(n-1)/2}$, which is interesting in its own right.
@@ 1 Introduction
The arrangement $\mathcal{A}(S)$ defined by a set $S$ of objects in the plane is the subdivision of the plane into faces, edges, and vertices induced by (the boundaries of) the objects. Arrangements in the plane and in higher dimensions play a fundamental role in computational geometry, since many geometric problems can be transformed into problems on arrangements. As a result, the combinatorial complexity of arrangements—the total number of vertices, edges, and faces (and higher-dimensional features)—induced by various types of objects has been studied extensively. These studies concern the complexity of complete arrangements as well as the complexity of various substructures: single cells, many cells, zones, levels, and so on; see the book by Sharir and Agarwal [7] or the surveys by Agarwal and Sharir [1] or Halperin and Sharir [5]. The known results on arrangements typically deal with constant-complexity objects such as lines, circles, or other objects whose boundaries intersect a bounded number of times. This is not surprising, as the complexity of the arrangement induced by some set of objects obviously cannot be bounded if we do not put a bound on the number of intersections between the boundaries of any two objects in the set. In this paper we are interested in whether anything meaningful can be said about complexity of an arrangement if the number of boundary intersections of any two objects in the arrangement is not necessarily bounded.
@@
Let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$.
Let $\Delta_{ij}$ denote the number of connected components of the intersection $D_{i}\cap D_{j}$ of a pair of objects $D_{i},D_{j}\in\mathcal{D}$, and define $\Delta:=\max_{i,j}\Delta_{ij}$. We refer to $\Delta$ as the overlap number of $\mathcal{D}$.
Note that the complexity of $\mathcal{A}$ cannot be bounded as a function of $n$ and $\Delta$. Indeed, even for $n=2$ and $\Delta=1$, the number of faces in $\mathcal{A}$ can be arbitrarily large, as illustrated in Figure 1(i).
@@
The second quantity of interest is the number of maximal faces in $\mathcal{A}$. Let ply of a face in $\mathcal{A}$ be the number of disks in $\mathcal{D}$ that contain the face. Two faces are called neighbors in $\mathcal{A}$ if they have at least one common edge, i.e., if their corresponding nodes in $\mathcal{G}^{*}$ are adjacent. A face is called maximal if its ply is larger than the ply of any of its neighboring faces.
@@
Let $\operatorname{diam}(\mathcal{G})$ denote the diameter of a graph $\mathcal{G}$, and let $\mu(\mathcal{A})$ denote the number of maximal faces of an arrangement $\mathcal{A}$. Furthermore, define
@@
$\mu(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{\mu(\mathcal{A}(\mathcal{D}))\},$
with the maximum taken over all sets $\mathcal{D}$ of $n$ topological disks with overlap number $\Delta$ in general position. We first consider the case of two disks. Bounding $\mu(\mathcal{A})$ is trivial for $n=2$, since $\mu(\mathcal{A})=\Delta$ if $D_{0}$ and $D_{1}$ intersect, and $\mu(\mathcal{A})=2$ if $D_{0}$ and $D_{1}$ are disjoint. Thus, $\mu(2,\Delta)=\max\{2,\Delta\}$. Bounding $\xi(2,\Delta)$ is less straightforward. We study this problem in Section 2, where we prove the following tight bound on the maximum possible diameter of $\mathcal{G}^{*}(\mathcal{A}(\mathcal{D}))$ in terms of $\Delta$.
@@
In Section 3 we then turn our attention to the much more challenging case of general $n$. Here, we first prove that $\mu(n,\Delta)$ is bounded.
@@
For any $n>2$, we have $\mu(n,\Delta)\leqslant n(\Delta+1)^{n(n-1)/2}$.
@@
For any $n>2$, we have $\xi(n,\Delta)\leqslant 2n\cdot\min\{n,\Delta+1\}\cdot(\Delta+1)^{n(n-1)/2}$.
@@
The problem of bounding the diameter of $\mathcal{G}^{*}$ is equivalent to the following question. For two points $s,t\in\mathbb{R}^{2}$ not lying on the boundary of any disk in $\mathcal{D}$, let $d(s,t)$ be the minimum number of times one needs to cross a disk boundary when going from $s$ to $t$. Then, the question is: what is the maximum value of $d(s,t)$ over all such pairs $s$ and $t$? This is related to the concept of barrier coverage in sensor networks, which aims to ensure that all possible paths through a surveillance domain will intersect the boundary of a sensor region at least once. Cardei and Wu [4] and Meguerdichian et al. [6] provide comprehensive overviews of work on this topic. The notion of barrier coverage closest to our work is the one introduced by Bereg and Kirkpatrick [3], who define the thickness of a barrier as the minimum number of sensor region intersections of a path between any two regions.
A substantial difference between the paper by Bereg and Kirkpatrick and our paper is the way of counting the number of crossings of an $st$-path with the region boundaries: they count a boundary crossing only if the curve enters a region for the first time, while we count every boundary crossing. Cabello et al. [2] study the related problem of computing, for a given set of line segments and two points $s$ and $t$, the minimum number of segments crossed by an $st$-path, but they also count each segment at most once.
@@ 2 The case of two disks
First, notice that a colored path from $u$ to $v$ always exists since $u$ and $v$ are purple. Indeed, the corresponding faces $f_{u}$ and $f_{v}$ lie in $D_{0}\cap D_{1}$, so there is a curve $\gamma\subset D_{0}$ connecting $f_{u}$ and $f_{v}$ inside $D_{0}$. This curve corresponds to a colored path in $\mathcal{G}^{*}$. Let $\pi(u,v)$ be a shortest colored path from $u$ to $v$ in $\mathcal{G}^{*}$. Since $\pi(u,v)$ is a shortest path, any node occurs at most once on it. Moreover, $\pi(u,v)$ does not have white nodes, so it alternates between purple nodes and red or blue nodes. Since $\pi(u,v)$ starts and ends at a purple node and it contains at most $\Delta$ purple nodes, $\pi(u,v)$ has at most $2\Delta-2$ edges.
@@
The dual graph $\mathcal{G}^{*}$ of two disjoint disks has diameter 2. Therefore, to prove the lower bound, it suffices to construct an arrangement of two disks with the overlap number $\Delta\geqslant 2$ whose dual graph has diameter $2\Delta$. Fig. 4 exhibits such an arrangement for $\Delta=6$. Each spiral consists of $4k$ thin corridors, two vertical and two horizontal. In the figure $k=3$, but the example easily extends to larger $k$. Note that when increasing $k$ by 1, we increase $\Delta$ by 2 as each vertical corridor of the blue disk intersects a horizontal corridor of the red disk. Thus, in general, $k=\Delta/2$. In the arrangement, we can identify $2k-1=\Delta-1$ rings, each consisting of a red and a blue face at equal distance from the innermost blue face; see the right image in Fig. 4, where the rings are shown in black and gray. Going from the innermost blue face to the innermost ring requires two hops, going from a ring to the next ring requires two hops, and going from the outermost ring to the outermost blue face requires two hops. Thus, the distance between the innermost and outermost blue face is $2\Delta$. This construction gives an arrangement of two disks with overlap number $\Delta>1$ whose dual graph has diameter $2\Delta$, for even $\Delta\geqslant 2$. By omitting the last two corridors of both disks, we obtain such an arrangement for odd $\Delta>2$ as well.
@@ 3 The general case of $n$ disks
We now consider the case of general $n$, which turns out to be significantly harder than the case of two disks. The main difficulty lies in bounding the number of maximal faces; for two disks this number is simply the overlap number $\Delta$ (assuming the disks are not disjoint), but for general $n$ it is not even clear that the number of maximal faces is bounded. Our main effort lies in proving that this is the case (Section 3.1). Once we have shown that, establishing a bound on the diameter of the dual graph of the arrangement of disks is relatively easy (Section 3.2).

### v2

@@ Abstract
Let $\mathcal{D}=\{D_{0},\ldots,D_{n-1}\}$ be a set of $n$ topological disks in the plane and let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$. For two disks $D_{i},D_{j}\in\mathcal{D}$, let $\Delta_{ij}$ be the number of connected components of $D_{i}\cap D_{j}$, and let $\Delta:=\max_{i,j}\Delta_{ij}$. We show that the diameter of $\mathcal{G}^{*}$, the dual graph of $\mathcal{A}$, can be bounded as a function of $n$ and $\Delta$. Thus, any two points in the plane can be connected by a Jordan curve that crosses the disk boundaries a number of times bounded by a function of $n$ and $\Delta$. In particular, for the case of two disks, we prove that the diameter of $\mathcal{G}^{*}$ is at most $\max\{2,2\Delta\}$ and this bound is tight. For the general case of $n>2$ disks, we show that the diameter of $\mathcal{G}^{*}$ is $O(n^{3}2^{n}\Delta)$. We achieve this by proving that the number of maximal faces in $\mathcal{A}$—faces whose ply is more than the ply of their neighboring faces—is $O(n^{2}2^{n}\Delta)$. To this end, we first show that the number of maximum faces—faces whose ply is nn—is $O(n^{2}\Delta)$; the latter bound, which is of independent interest, is tight in the worst case.
@@ 1 Introduction
The arrangement $\mathcal{A}(S)$ defined by a set $S$ of objects in the plane is the subdivision of the plane into faces, edges, and vertices induced by (the boundaries of) the objects. Arrangements in the plane and in higher dimensions play a fundamental role in computational geometry, since many geometric problems can be transformed into problems on arrangements. As a result, the combinatorial complexity of arrangements—the total number of vertices, edges, and faces (and higher-dimensional features)—induced by various types of objects has been studied extensively. These studies concern the complexity of complete arrangements as well as the complexity of various substructures: single cells, many cells, zones, levels, and so on; see the book by Sharir and Agarwal [DBLP:books/daglib/0080837] or the surveys by Agarwal and Sharir [DBLP:books/el/00/AgarwalS00a] or Halperin and Sharir [DBLP:reference/cg/arrangements]. The known results on arrangements typically deal with constant-complexity objects such as lines, circles, or other objects whose boundaries intersect a bounded number of times. This is not surprising, as the complexity of the arrangement induced by some set of objects obviously cannot be bounded if we do not put a bound on the number of intersections between the boundaries of any two objects in the set. In this paper we are interested in whether anything meaningful can be said about complexity of an arrangement if the number of boundary intersections of any two objects in the arrangement is not necessarily bounded.
@@
Let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$. Let $\Delta_{ij}$ denote the number of connected components of the intersection $D_{i}\cap D_{j}$ of a pair of objects $D_{i},D_{j}\in\mathcal{D}$, and define $\Delta:=\max_{i,j}\Delta_{ij}$. We refer to $\Delta$ as the overlap number of $\mathcal{D}$. Note that the complexity of $\mathcal{A}$ cannot be bounded as a function of $n$ and $\Delta$. Indeed, even for $n=2$ and $\Delta=1$, the number of faces in $\mathcal{A}$ can be arbitrarily large, as illustrated in Figure 1(i).
@@
The second quantity of interest is the number of maximal faces in $\mathcal{A}$. Let ply of a face in $\mathcal{A}$ be the number of disks in $\mathcal{D}$ that contain the face. Two faces are called neighbors in $\mathcal{A}$ if they have at least one common edge, if their corresponding nodes in $\mathcal{G}^{*}$ are adjacent. A face is called maximal if its ply is larger than the ply of any of its neighboring faces. A special case of a maximal face is a maximum face, which is a face whose ply is $n$ or, in other words, a face that is contained in all disks in $\mathcal{D}$.
@@
Let $\operatorname{diam}(\mathcal{G})$ denote the diameter of a graph $\mathcal{G}$, and let $\mu(\mathcal{A})$ and $M(\mathcal{A})$ denote the number of maximal and maximum faces, respectively, of the arrangement $\mathcal{A}$. Furthermore, define
@@
$\mu(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{\mu(\mathcal{A}(\mathcal{D}))\}$
and
$M(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{M(\mathcal{A}(\mathcal{D}))\},$
with the maximum taken over all sets $\mathcal{D}$ of $n$ topological disks with overlap number $\Delta$ in general position. We first consider the case of two disks. Bounding $\mu(\mathcal{A})$ is trivial for $n=2$, since $\mu(\mathcal{A})=\Delta$ if $D_{0}$ and $D_{1}$ intersect, and $\mu(\mathcal{A})=2$ if $D_{0}$ and $D_{1}$ are disjoint. Thus, $\mu(2,\Delta)=\max\{2,\Delta\}$ and $M(2,\Delta)=\Delta$. Bounding $\xi(2,\Delta)$ is less straightforward. We study this problem in Section 2, where we prove the following tight bound on the maximum possible diameter of $\mathcal{G}^{*}(\mathcal{A}(\mathcal{D}))$ in terms of $\Delta$.
@@
In Section 3 we then turn our attention to the much more challenging case of general $n$. Here, we first prove the following upper bound on the number of maximum faces.
@@
For any $n>2$, we have $M(n,\Delta)<2n(n-1)\Delta$.
We also give a construction showing that $M(n,\Delta)=\Omega(n^{2}\Delta)$, showing that the bound from Theorem 1.2 is asymptotically tight. This result implies the following bound on the number of maximal faces.
Theorem 1.3.
For any $n>2$, we have $\mu(n,\Delta)=O(n^{2}2^{n}\Delta)$.
@@
For any $n>2$, we have $\xi(n,\Delta)=O(n^{3}2^{n}\Delta)$.
@@
The problem of bounding the diameter of $\mathcal{G}^{*}$ is equivalent to the following question. For two points $s,t\in\mathbb{R}^{2}$ not lying on the boundary of any disk in $\mathcal{D}$, let $d(s,t)$ be the minimum number of times one needs to cross a disk boundary when going from $s$ to $t$. Then, the question is: what is the maximum value of $d(s,t)$ over all such pairs $s$ and $t$? This is related to the concept of barrier coverage in sensor networks, which aims to ensure that all possible paths through a surveillance domain will intersect the boundary of a sensor region at least once. Cardei and Wu [cardei2004coverage] and Meguerdichian et al. [meguerdichian2001coverage] provide comprehensive overviews of work on this topic. The notion of barrier coverage closest to our work is the one introduced by Bereg and Kirkpatrick [bereg2009approximating], who define the thickness of a barrier as the minimum number of sensor region intersections of a path between any two regions. A substantial difference between the paper by Bereg and Kirkpatrick and our paper is the way of counting the number of crossings of an $st$-path with the region boundaries: they count a boundary crossing only if the curve enters a region for the first time, while we count every boundary crossing. Cabello et al. [DBLP:journals/ijcga/AltCGK17] study the related problem of computing, for a given set of line segments and two points $s$ and $t$, the minimum number of segments crossed by an $st$-path, but they also count each segment at most once.
@@ 2 The case of two disks
First, notice that a colored path from $u$ to $v$ always exists since $u$ and $v$ are purple. Indeed, the corresponding faces $f_{u}$ and $f_{v}$ lie in $D_{0}\cap D_{1}$, so there is a curve $\gamma\subset D_{0}$ connecting $f_{u}$ and $f_{v}$ inside $D_{0}$. This curve corresponds to a colored path in $\mathcal{G}^{*}$.
Let $\pi(u,v)$ be a shortest colored path from $u$ to $v$ in $\mathcal{G}^{*}$. Since $\pi(u,v)$ is a shortest path, any node occurs at most once on it. Moreover, $\pi(u,v)$ does not have white nodes, so it alternates between purple nodes and red or blue nodes. Since $\pi(u,v)$ starts and ends at a purple node and it contains at most $\Delta$ purple nodes, $\pi(u,v)$ has at most $2\Delta-2$ edges.
@@
The dual graph $\mathcal{G}^{*}$ of two disjoint disks has diameter 2. Therefore, to prove the lower bound, it suffices to construct an arrangement of two disks with the overlap number $\Delta\geqslant 2$ whose dual graph has diameter $2\Delta$. Fig. 4 exhibits such an arrangement for $\Delta=6$. Each spiral consists of $4k$ thin corridors, two vertical and two horizontal. In the figure $k=3$, but the example easily extends to larger $k$. Note that when increasing $k$ by 1, we increase $\Delta$ by 2 as each vertical corridor of the blue disk intersects a horizontal corridor of the red disk. Thus, in general, $k=\Delta/2$. In the arrangement, we can identify $2k-1=\Delta-1$ rings, each consisting of a red and a blue face at equal distance from the innermost blue face; see the right image in Fig. 4, where the rings are shown in black and gray. Going from the innermost blue face to the innermost ring requires two hops, going from a ring to the next ring requires two hops, and going from the outermost ring to the outermost blue face requires two hops. Thus, the distance between the innermost and outermost blue face is $2\Delta$.
This construction gives an arrangement of two disks with overlap number $\Delta>1$ whose dual graph has diameter $2\Delta$, for even $\Delta\geqslant 2$. By omitting the last two corridors of both disks, we obtain such an arrangement for odd $\Delta>2$ as well.
@@ 3 The general case of $n$ disks
We now consider the case of general $n$, which turns out to be significantly harder than the case of two disks. The main difficulty lies in bounding the number of maximal faces; for two disks this number is simply the overlap number $\Delta$ (assuming the disks are not disjoint), but for general $n$ it is not even clear that the number of maximal faces (or the number of maximum faces, for that matter) is bounded. Our main effort lies in proving that this is the case (Section 3.1). Once we have shown that, establishing a bound on the diameter of the dual graph of the arrangement of disks is relatively easy (Section 3.2).
3.1 The proof of Theorems 1.2 and 1.3
Recall that $\mu(A)$ and $M(\mathcal{A})$ denote the number of maximal and maximum faces, respectively, in the arrangement $\mathcal{A}$ defined by the set $\mathcal{D}$ of $n$ disks. We first prove our bound on the number of maximum faces. See 1.2
Proof 3.1.
Clearly, $M(2,\Delta)=\Delta$ by definition. We will show that for $n\geqslant 3$ we have
$M(n,\Delta)\leqslant M(n-1,\Delta)+4(n-1)(\Delta-1).$ (1)
This implies the theorem, since
$\begin{array}[]{lll}M(n,\Delta)&\leqslant&M(2,\Delta)+\sum_{j=3}^{n}4(j-1)(\Delta-1)\\ &=&\Delta+\sum\limits_{j=3}^{n}4(j-1)(\Delta-1)\\ &<&\sum\limits_{j=2}^{n}4(j-1)\Delta\mbox{ }\\ &=&2n(n-1)\Delta.\end{array}$
To prove that Inequality (1) holds, consider a set $\mathcal{D}$ of $n\geqslant 3$ disks with overlap number $\Delta$, and let $D_{0}$ be an arbitrary disk in $\mathcal{D}$. Define $\mathcal{D}_{1}:=\mathcal{D}\setminus\{D_{0}\}$ and let $D_{1},\ldots,D_{n-1}$ be the disks in $\mathcal{D}_{1}$. We will show that
$M(\mathcal{A}(\mathcal{D}))\leqslant M(\mathcal{A}(\mathcal{D}_{1}))+4(n-1)(\Delta-1).$
Note that some faces do not change when we remove $D_{0}$ from $\mathcal{D}$, while other faces—the ones incident to $\partial D_{0}$—expand into larger faces. Now consider a maximum face $f$ of $\mathcal{A}(\mathcal{D})$ and let $\mathit{parent}(f)$ be the face in $\mathcal{A}(\mathcal{D}_{1})$ that contains $f$. Note that $\mathit{parent}(f)$ is a maximum face in $\mathcal{A}(\mathcal{D}_{1})$. We consider two types of maximum faces in $\mathcal{A}(\mathcal{D})$, illustrated in Figure 5(i).
Figure 5: (i) An arrangement of three disks. Faces $f_{1}$ and $f_{4}$ are safe, faces $f_{2}$ and $f_{3}$ are unsafe. (ii) The red disk defines four dangerous intervals along $\partial D_{0}$, indicated by thick lines.
  • A safe face is a maximum face $f$ in $\mathcal{A}(\mathcal{D})$ such that $f$ is the unique maximum face in $\mathcal{A}(\mathcal{D})$ contained in $\mathit{parent}(f)$.
  • An unsafe face is a maximum face $f$ in $\mathcal{A}(\mathcal{D})$ such that there is a maximum face $f^{\prime}\neq f$ in $\mathcal{A}(\mathcal{D})$ that is also contained in $\mathit{parent}(f)$.
The number of safe faces is obviously at most $M(\mathcal{D}_{1})$, so
[…]
Note that the inequality is strict, unless there are no unsafe faces. Indeed, any set of unsafe faces with the same parent face already contributes 1 to $M(\mathcal{D}_{1})$. Next, we bound the number of unsafe faces.
Consider the arrangement $\mathcal{A}(\{D_{0},D_{i}\})$ induced by $D_{0}$ and $D_{i}$. Any vertex $v$ of this arrangement is incident to two edges of $\mathcal{A}(\{D_{0},D_{i}\})$ contained in $\partial D_{i}$, one inside $D_{0}$ and one outside $D_{0}$. We call the latter edge the outside curve of $v$ and we denote it by $\gamma(v)$. Let $\mathcal{I}(D_{i}):=D_{i}\cap\partial D_{0}$ be the (possibly empty) set of intervals on $\partial D_{0}$ covered by $D_{i}$. Consider an interval $I\in\mathcal{I}(D_{i})$ and let $v$ and $w$ be its endpoints. We call the interval $I$ dangerous if $\gamma(v)\neq\gamma(w)$; see Figure 5(ii) for an example.
Claim 1.
Any disk $D_{i}\in\mathcal{D}_{1}$ defines at most $2\Delta-2$ dangerous intervals.
{claimproof}
Assume there is at least one dangerous interval, otherwise the claim holds trivially. Let $F_{i}$ be the set of faces of $\mathcal{A}(\{D_{0},D_{i}\})$ that are incident to a dangerous interval, let $F_{i}^{\mathrm{in}}\subset F_{i}$ be the faces that lie inside $D_{0}$ and let $F_{i}^{\mathrm{out}}\subset F_{i}$ be the faces that lie outside $D_{0}$. Let $\mathcal{G}$ be the graph whose nodes are the faces in $F_{i}$ and where there is an edge between two faces if they share a dangerous interval; see Figure 6.
Figure 6: The graph $\mathcal{G}$ used in the proof of 1.
$\mathcal{G}$ is a bipartite graph with node sets $F_{i}^{\mathrm{in}}$ and $F_{i}^{\mathrm{out}}$, and that $|F_{i}^{\mathrm{in}}|\leqslant\Delta$ by definition of the overlap number. Moreover, any node $f\in F_{i}^{\mathrm{out}}$ has degree at least two in $\mathcal{G}$. Indeed, if $f$ had degree one, then the face $f$ would be adjacent to a single dangerous interval, which is impossible by the definition of dangerous intervals. Note that $\mathcal{G}$ is a tree, since $D_{i}$ is a topological disk.
We now view $\mathcal{G}$ as a rooted tree, where we pick an arbitrary node $f\in F_{i}^{\mathrm{out}}$ as the root. Since the nodes in $F_{i}^{\mathrm{out}}$ have degree at least 2 in $\mathcal{G}$, they are not leaves. To bound $|F_{i}^{\mathrm{out}}|$, we charge each node in $F_{i}^{\mathrm{out}}$ to an arbitrary one of its children. Thus we charge each node in $F_{i}^{\mathrm{out}}$ to a node in $F_{i}^{\mathrm{in}}$ in such a way that each node in $F_{i}^{\mathrm{in}}$ is charged at most once. Since the root node has at least two children, one of which will not be charged, we conclude that
$|F_{i}^{\mathrm{out}}|\leqslant|F_{i}^{\mathrm{in}}|-1\leqslant\Delta-1.$
Hence, the total number of nodes in $\mathcal{G}$ is at most $2\Delta-1$, and so the number of edges in $\mathcal{G}$, which equals the number of dangerous intervals, is at most $2\Delta-2$. We call an edge of a face $f$ of $\mathcal{A}(\mathcal{D})$ a boundary edge of $f$ if it is contained in $\partial D_{0}$. The next claim relates the number of dangerous intervals to the number of unsafe faces.
Claim 2.
Let $f$ be an unsafe face. Then $f$ has a boundary edge $e$ such that at least one of the endpoints of $e$ is an endpoint of a dangerous interval.
{claimproof}
Let $f^{\prime}\neq f$ be a maximum face in $\mathcal{A}(\mathcal{D})$ that is contained in $\mathit{parent}(f)$; such a face exists because $f$ is unsafe. Then there is a curve $\xi$ in the interior of $\mathit{parent}(f)$ that connects $f$ to $f^{\prime}$. (More precisely, $\xi$ connects a point in the interior of $f$ to a point in the interior of $f^{\prime}$.) Since $f$ and $f^{\prime}$ are distinct faces in $\mathcal{A}(\mathcal{D})$, this curve $\xi$ must cross a boundary edge, $e$, of $f$. We can assume without loss of generality that $\xi$ crosses $e$ once. Let $v$ and $w$ be the endpoints of $e$, and assume for a contradiction that neither $v$ nor $w$ is an endpoint of a dangerous interval. Let $\gamma(v)$ and $\gamma(w)$ be the outside curves of $v$ and $w$, respectively, and let $D_{i}$ and $D_{j}$ be such that $\gamma(v)\subset\partial D_{i}$ and $\gamma(w)\subset\partial D_{j}$ We consider two cases.
Figure 7: Illustration for the proof of 2.
  • If $i=j$ then $\gamma(v)=\gamma(w)$, because we assumed that neither $v$ nor $w$ is an endpoint of a dangerous interval. Hence, $f^{\prime}$ is contained in the region $R$ enclosed by $\gamma(v)\cup e$; see Figure 7(i). But $R$ lies fully outside $D_{0}$, which contradicts that $f^{\prime}$ is a maximum face in $\mathcal{A}(\mathcal{D})$.
  • If $i\neq j$ then $\gamma(v)$ must intersect $\gamma(w)$. To see this, consider the interval $I_{v}\subset\partial D_{0}\cap D_{i}$ with endpoint $v$ and containing $e$, and the interval $I_{w}\subset\partial D_{0}\cap D_{j}$ with endpoint $w$ and containing $e$. Then $w\in I_{v}$ and $v\in I_{w}$. Together with the fact that $I_{v}$ and $I_{w}$ are not dangerous by assumption, this implies that $\gamma(v)$ indeed intersects $\gamma(w)$; see Figure 7(ii). Let $R$ be the bounded face in the arrangement defined by $e$, $\gamma(v)$ and $\gamma(w)$ that is incident to $e$. Then $f^{\prime}$ is contained in $R$, which lies outside $D_{0}$, contradicting that $f^{\prime}$ is a maximum face in $\mathcal{A}(\mathcal{D})$.
We derived a contradiction in both cases, thus finishing the proof of the claim. 2 allows us to charge each unsafe face to a dangerous interval in such a way that each unsafe interval is charged at most once for each of its endpoints. By 1, the total number of dangerous intervals is at most $(n-1)(2\Delta-2)$. Hence, the number of unsafe faces is at most $2(n-1)(2\Delta-2)$. This finishes the proof of Inequality (1) and, hence, of the theorem.
The construction in Figure 8 shows that $M(n,\Delta)=\Omega(n^{2}\Delta)$.
Figure 8: A construction showing that $M(n,\Delta)=\Omega(n^{2}\Delta)$.
Theorem 1.2 implies the following bound on the number of maximal faces.
See 1.3
Proof 3.2.
We claim that for any $n\geqslant 2$ we have $\mu(n,\Delta)\leqslant(2^{n}-1)\cdot M(n,\Delta)$. Together with Theorem 1.2, this implies the theorem.
To prove the claim, let $\mathcal{D}$ be a collection of $n$ topological disks with overlap number $\Delta$, and let $f$ be a maximal face of $\mathcal{A}(\mathcal{D})$. Then $f$ is a maximum face in the arrangement $\mathcal{A}(\mathcal{D}_{f})$, where $\mathcal{D}_{f}\subseteq\mathcal{D}$ is the set of disks containing $f$. There are at most $M(|\mathcal{D}_{f}|,\Delta)$ such faces, which we charge to $D_{f}$. Hence,
$\mu(n,\Delta)\leqslant\sum_{\emptyset\neq\mathcal{D}^{\prime}\subseteq\mathcal{D}}M(|\mathcal{D}^{\prime}|,\Delta)\leqslant\sum_{\emptyset\neq\mathcal{D}^{\prime}\subseteq\mathcal{D}}M(n,\Delta)=(2^{n}-1)\cdot M(n,\Delta),$
as claimed. For the second inequality, note that $M(n,\Delta)$ is non-decreasing in $n$.